OpenStudy (fatimab2000):

Greg has only nickels and dimes in his money box. He knows that he has more than $10 in the box. Let x represent the number of nickels in the box and y represent the number of dimes in the box. Which of the following statements best describes the steps to graph the solution to the inequality in x and y? Draw a dashed line to represent the graph of 5x + 10y = 1000 and shade the portion below the line for positive values of x and y. Draw a dashed line to represent the graph of 5x + 10y = 1000 and shade the portion above the line for positive values of x and y.

OpenStudy (fatimab2000):

Draw a dashed line to represent the graph of 10x + 5y = 1000 and shade the portion above the line for positive values of x and y. Draw a dashed line to represent the graph of 10x + 5y = 1000 and shade the portion below the line for positive values of x and y.

OpenStudy (fatimab2000):

@ganeshie8

OpenStudy (fatimab2000):

will give medal !!!!!

ganeshie8 (ganeshie8):

`Greg has only nickels and dimes in his money box. He knows that he has more than $10 in the box.` if x is the number of nickels y is the number of dimes then, do we get the inequality \[5x+10y\gt 1000\] ?

ganeshie8 (ganeshie8):

notice that i have converted nickels, dimes, and dollars into cents

OpenStudy (fatimab2000):

i dont get it

ganeshie8 (ganeshie8):

how many cents make a nickel ?

OpenStudy (fatimab2000):

5

OpenStudy (fatimab2000):

so 5x is the nickels

ganeshie8 (ganeshie8):

yes, having x nickels is same as having 5x cents

ganeshie8 (ganeshie8):

how about dimes ? how many cents make a dime ?

OpenStudy (fatimab2000):

10

ganeshie8 (ganeshie8):

so, having y dimes is same as having 10y cents

ganeshie8 (ganeshie8):

how about dollars ? how many cents make a dollar ?

OpenStudy (fatimab2000):

100

ganeshie8 (ganeshie8):

good, now go through my first reply

OpenStudy (fatimab2000):

10x + 5y = 1000

OpenStudy (fatimab2000):

=100

ganeshie8 (ganeshie8):

10$ = 10*100 cents = 1000 cents

OpenStudy (fatimab2000):

the answer would be c right

OpenStudy (fatimab2000):

@ganeshie8

ganeshie8 (ganeshie8):

nope, read my first reply again

OpenStudy (fatimab2000):

sorry it would be B

ganeshie8 (ganeshie8):

Yes Draw a dashed line to represent the graph of 5x + 10y = 1000 and shade the portion above the line for positive values of x and y.

OpenStudy (fatimab2000):

thx